Find the time taken by sunlight to reach the ground if the distance between the sun and earth is 1.5x10 8 km, velocity of light is 3x10 8 m/s.
1
Expert's answer
2020-12-06T17:31:25-0500

The time "t" may be obtained as

"t = \dfrac{S}{v}" , where "S" is the distance between the Sun and Earth and "v" is the velocity of light. Therefore,

"t = \dfrac{1.5\cdot10^8\cdot10^3\,\mathrm{m}}{3\cdot10^8\,\mathrm{m/s}} = 500\,\mathrm{s} = 8.3\,\mathrm{min}."